# Mitigating Nonlinear Systematics in Weak Lensing Surveys II: Stability and Diagnostics with Intrinsic Alignment

### v1

@@ Abstract
The Bernardeau–Nishimichi–Taruya (BNT) transform provides a powerful framework for analysing tomographic cosmic shear data by improving the localization of shear correlations in physical scale. It operates by performing a linear combination of the shear data vector in $\ell$-space, yielding a transformed vector that is better localized in both redshift and $k$-space. BNT is particularly useful for estimating cosmological parameters while minimizing the impact of poorly understood nonlinear physics, without discarding large amounts of information as is typically done with simple scale cuts. In our previous work, we showed that BNT outperforms traditional weak-lensing analyses; however, that study did not include intrinsic alignments (IA). In the present work, we assess the robustness of our BNT-based $k$-cut framework in the presence of realistic IA models. We consider two cases: (i) when the assumed IA model used in sampling is close to, but not identical to, the true one, and (ii) when the assumed IA model is significantly mis-specified. In the first case, the $k$-cut framework yields precise and unbiased $S_{8}$ constraints even with limited knowledge of large-scale modes. In the second, the BNT transform serves as a powerful diagnostic tool, revealing internal inconsistencies in the assumed IA model across $k$- and redshift-space, indicating that the corresponding cosmological constraints cannot be trusted until a more accurate IA model is adopted.
@@ I Introduction
Weak gravitational lensing is fundamentally a two-dimensional probe: for a given source redshift, it measures anisotropies in the mass distribution projected along the line of sight from the sources to the observer. This makes it challenging to identify biases in the nonlinear three-dimensional modeling of the matter power spectrum, because only two-dimensional projections are observed. Consequently, the usual approach is to apply scale cuts to the WL data vector, at the cost of significant information loss. However, [4] showed that a certain linear combination of weak-lensing observables can recover some of the three-dimensional information. This is possible because, for a lens at a given redshift, the redshift scaling of the WL effect is purely geometric and independent of the matter distribution—a property that underlies nulling techniques [5, 6]. The specific linear combination that achieves this, developed in [4], is called the Bernardeau–Nishimichi–Taruya (BNT) transform. In our previous work [7] (Paper I), we adapted and tested this transform as a targeted solution for mitigating nonlinear systematics.
We showed that a BNT-transformed data vector, combined with a well-defined angular scale filtering ($\ell$-cut in harmonic space), can effectively operate as a $k$-cut, enabling more robust handling of uncertainties in the nonlinear modeling of the three-dimensional matter power spectrum. However, in Paper I we did not account for intrinsic alignment (IA), a known contaminant of the WL signal that can mimic lensing by correlating galaxy shapes over large distances. In this paper, we revisit a number of results obtained in paper I and reevaluate the performance of the BNT transform when IA is taken into account.
@@ II.1 Cosmic Shear
We consider a cosmological survey with measured photometric redshifts for galaxies, enabling tomographic binning with source-galaxy distributions $n_{i}(\chi)$ for the $i$th bin. The two-dimensional cosmic-shear power spectrum $C^{i,j}(\ell)$ between tomographic bins $(i,j)$ is given by the line-of-sight projection of the non-linear three-dimensional matter power spectrum $P_{\rm nl}(k;z)$, as follows [8]:
@@
$W_{\gamma}^{i}(\chi)=\dfrac{\Omega_{m}^{2}H_{0}^{4}}{c^{2}}\int\text{d}\chi^{\prime}\dfrac{n_{i}(\chi^{\prime})}{a(\chi)}\dfrac{(\chi^{\prime}-\chi)\chi}{\chi^{\prime}},$ (2)
where $a(\chi)$ is the scale factor. We have assumed a flat Universe $K=0$, for non-flat geometry, the radial comoving distance $\chi$ has to be replaced by the comoving angular diameter distance $f_{K}(\chi)$ in Eq. 2:
@@ II.2 Intrinsic Alignment
In addition to the coherent distortions induced by gravitational lensing described by Eq. 1, intrinsic alignments (IA) generate additional angular correlations among galaxy shapes—both within a given tomographic bin, through local tidal interactions, and across widely separated bins, via long-range correlations. These intrinsic correlations can mimic or bias the lensing signal, making IA a significant astrophysical systematic in weak-lensing analyses [9].
@@
where $\rho_{\mathrm{crit}}$ is the critical density of the universe, $D(z)$ is the structure growth function, $z_{0}$ is a pivot redshift often chose to be $z_{0}=0.62$ [12], and $\tilde{C}_{1}=5\times 10^{-14}\,h^{-2}\mathrm{M}_{\odot}^{-1}\mathrm{Mpc}^{3}$ is a normalisation constant.
@@
where $\langle a(z)\rangle_{i}$ is the $n(z)$-weighted average scale factor in tomographic bin $i$, and $a_{\mathrm{piv}}=0.769$ is the pivotal scale factor. Note that, unlike the $\eta_{\mathrm{TA}}$ parameterisation of NLA model, this NLA-z model can change sign as a function of redshift.
@@
$\gamma_{ij}^{\mathrm{IA}}=C_{\mathrm{1}}s_{ij}+C_{1\delta}(\delta\times s_{ij})+C_{2}\left[\sum_{k=0}^{2}s_{ik}s_{kj}-\frac{1}{3}\delta_{ij}s^{2}\right]+$... (6)
@@
where $A_{\mathrm{TT}}$ is the Tidal Torquing scaling amplitude and $b_{\mathrm{TA}}>0$ is the density weighting bias. To avoid introducing too many equations, please refer to the Section III. of [14] for the perturbation theory-based prescription to compute $P_{\mathrm{GI}}(k)$ and $P_{\mathrm{II}}(k)$ under the TATT model. It is worth mentioning that, for the TATT model, Eq. 3 cannot be applied. This is a consequence of Eq. 6 which implies that the redshift and $k$ dependence can no longer be factorized.
After obtaining the power spectra $P_{\mathrm{GG}}$, $P_{\mathrm{GI}}$ and $P_{\mathrm{II}}$, we can project it to the 2-D celestial sphere with a similar procedure mentioned in the Eq. 1:
$C_{\gamma^{i},\gamma^{j}}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{\gamma}(\chi)W^{j}_{\gamma}(\chi)P_{\mathrm{GG}}(k;z)$ (9)
$C_{\gamma^{i},I^{j}}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{\gamma}(\chi)W^{j}_{N}(\chi)P_{\mathrm{GI}}(k;z)$ (10)
$C_{I^{i},\gamma^{j}}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{N}(\chi)W^{j}_{\gamma}(\chi)P_{\mathrm{GI}}(k;z)$ (11)
$C_{I^{i},I^{j}}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{N}(\chi)W^{j}_{N}(\chi)P_{\mathrm{II}}(k;z)$ (12)
and
$C^{i,j}\equiv C_{\gamma^{i},\gamma^{j}}+C_{\gamma^{i},I^{j}}+C_{I^{i},\gamma^{j}}+C_{I^{i},I^{j}},$ (13)
where $W^{i}_{N}(\chi)$ is the number density projection kernel:
$W^{i}_{N}(\chi)=b^{i}\frac{H(\chi)}{c}\,n_{i}(z(\chi)),$ (14)
and $b^{i}$ being the linear galaxy bias of ith tomographic bin. In this paper, in order to keep the focus on the intrinsic alignment, we set $b^{i}=1$. For an effective description of the intrinsic alignment of its variation compared to the standard NLA model (Eq. 4), we also define an effective tidal alignment amplitude $A_{\mathrm{TA,eff}}$:
$A_{\mathrm{TA,eff}}(k,z)$ $\equiv$ $\dfrac{D(z)}{\tilde{C}_{1}\rho_{\mathrm{crit}}\Omega_{m}}\sqrt{\dfrac{P_{\mathrm{II}}(k,z)}{P_{\mathrm{GG}}(k,z)}}$
$=$ $-\dfrac{D(z)}{\tilde{C}_{1}\rho_{\mathrm{crit}}\Omega_{m}}C_{1}(z)$
@@ II.3 BNT Transform
It is worth mentioning that, for the NLA model, combining Eq. 3 and Eqs. 9-12 leads to a compact form for the BNT data vector $\hat{C}^{a,b}(\ell)$:
@@
In the rest of this work, we will use Euclid-like mock with ten redshift bins [16]. Figure 2 displays the three possible kernel combinations as functions of $\chi$, alongside the effective kernel $\chi^{-2}\hat{W}^{2}_{\mathrm{eff}}(\chi)$. The BNT intrinsic-alignment kernel is less localized in redshift than the BNT lensing kernel, yet it remains more localized than the original noBNT lensing kernel.
@@ III.1 Scale Cuts
where $k_{\mathrm{min}}=0.0025$ Mpc-1 and $k_{\mathrm{max}}=100$ Mpc-1 is the entire computational domain in $k$-space. As noted in Paper I, the BNT transform does not need to be computed with the true cosmology. Actually, searching for the true cosmology for which nulling would be exact can itself be exploited ([17, 18]).
Following Paper I, we also define a fractional difference threshold $\mathcal{T}_{\mathrm{FD}}$ to characterise how a $\ell$-cut is defined based on the tolerance region with $\hat{\mathfrak{R}}=1\pm\mathcal{T}_{\mathrm{FD}}$. In order to include the impact of both $k_{\mathrm{cuts}}$ and $\mathcal{T}_{\mathrm{FD}}$, we use the same continuous Boolean weight function $\mathcal{V}(k_{\mathrm{cuts}},\mathcal{T}_{\mathrm{FD}})$ as the paper I:
@@ IV.1 BNT and Intrinsic Alignment
We first examine the joint constraints on $(S_{8},\Omega_{m})$ in a setting where the fiducial nonlinear power spectrum is biased relative to the model used for sampling. As in Paper I, we compare constraints from the BNT and noBNT approaches for various choices of $k_{\rm cut}$, using AxionHMcode with non-zero intrinsic alignments for the fiducial data vector, while the sampling model uses HMcode. In Figure 4, we show the cosmological constraints for different physical scale cuts $k_{\rm cut}$, sampled over a TATT model with $b_{\mathrm{TA}}=0.0$, compared to the fiducial TATT model with $b_{\mathrm{TA}}=0.21$. Notably, this mismatch in $b_{\mathrm{TA}}$ does not introduce any significant additional bias in the cosmological constraints. The bias in $S_{8}$ and $\Omega_{m}$ appears when small-scale information is included, as shown by the shift of contours in Figure 4, which is primarily due to the impact of ultralight axions. However, as we progressively restrict the analysis to larger scales, the $S_{8}$ and $\Omega_{m}$ constraints converge toward the “true background cosmology” indicated by the cross in the contour plot. At $k_{\mathrm{cut}}=0.1\;\mathrm{Mpc}^{-1}$, the recovered cosmology aligns well with the fiducial. These results closely match the result in Paper I, where intrinsic alignment was not included. We conclude that the performance of BNT is not degraded by the presence of intrinsic alignment. By contrast, the noBNT constraints lose most of their constraining power in this regime, a result also similar to Paper I. These results suggest that, even with good but imperfect knowledge of the IA model, the BNT-transform-based $\ell$-cut method remains effective for delivering unbiased cosmological constraints.
@@ IV.2 Biases induced by Intrinsic Alignment Models
We first investigate the impact of an incorrect IA model in the redshift space by using the NLA-z model from Table 2 for fiducial and NLA model for the sampling. The NLA-z model introduces a different redshift dependence than the standard NLA model, which illustrates the impact of choosing an incorrect IA model. The data vector is split in three tomographic redshift groups: $(i,j)$ or $(a,b)$ for noBNT or BNT, belonging to either $\{1,2,3\}$, $\{4,5,6\}$, or $\{7,8,9,10\}$. For each subset, we exclude cross-correlations between bins belonging to different groups, as explained in Section III.2. Figure 6 shows the resulting constraints on $A_{\mathrm{TA}}$ and $\eta_{\mathrm{TA}}$ derived from the posterior sampling. In both the noBNT and BNT analyses, the resulting contours display substantial inconsistencies across different redshift-bin groupings. Constraints from the first three tomographic bins deviate from those obtained using the full redshift range by more than $5\sigma$ in $\eta_{\mathrm{TA}}$. Likewise, the constraints from the last four tomographic bins differ from those of the first three by over $5\sigma$, while still showing a tension exceeding $4\sigma$ relative to the unsliced case. These findings demonstrate that strong internal inconsistencies can arise when the IA model is incorrect. Redshift grouping therefore provides a useful diagnostic for testing the validity of the assumed IA model.
@@
Figure 7: The effective tidal alignment amplitude $A_{\mathrm{TA,eff}}(z)$ (Eq. II.2) constraints of all three redshift slices derived from the contour of the Figure 6, for both BNT and noBNT cases. The z-values and their errorbars of the data points are derived from the weighted mean and weighted standard deviation, when all galaxies within the $z$-slice (like $\mathbb{Z}_{1,2,3}$) are categorized to be within one single redshift distribution. For the illustration purpose only, the z-value of the noBNT and BNT data points are each shifted by $\pm 0.01$ on the plot, with their $A_{\mathrm{TA,eff}}(z)$ kept to its original constraint value.
@@
Based on the BNT and noBNT constraints shown in Figure 6, we fit an effective tidal alignment amplitude, $A_{\mathrm{TA,eff}}(z)$ (using the definition from Eq.II.2) for each redshift grouping. Figure 7 presents the resulting $A_{\mathrm{TA,eff}}(z)$ compared to the input NLA-z Fiducial model. The centre values and bin width along the $z$-axis are obtained from the $n(z)$-weighted mean and standard deviation, by combining the redshift slices $\mathbb{Z}_{1,2,3}$, $\mathbb{Z}_{4,5,6}$, and $\mathbb{Z}_{7,8,9,10}$ into three effective single redshift distributions. This demonstrates that, for an intrinsic alignment model with explicit redshift dependence, such an approach can serve as a viable method to constrain the detailed redshift evolution of the IA model if the sampling model cannot recover it. For instance, $A_{\mathrm{IA,eff}}(z)$ (Eq. II.2) cannot cross zero with our sampling IA model (Eq. 4), but it can do so with the fiducial IA model (Eq. 5), therefore creating a discrepancy.
@@
Similarly, we investigate the impact of $k$-band cuts when the fiducial and sampling IA models differ. In this test, we adopt the $k$-dependent nonlinear TATT model, which introduces additional clustering power at large $k$-scales compared to the conventional NLA model. The fiducial data vector is generated using the TATT model with parameters fitted from the Samuroff et al. [21] study, while sampling is performed with the standard NLA model. Figure 8 shows the constraints on ($\eta_{\rm TA},A_{\rm TA}$) for three $k$-bands: $k_{\mathrm{cuts}}=(0.033,0.33)$, $(0.1,1.0)$, and $(0.33,3.3)\;\mathrm{Mpc}^{-1}$ . Notably, the noBNT analysis shows a substantial loss of constraining power, particularly for the intrinsic-alignment parameter $A_{\mathrm{TA}}$. In contrast, the contours of the BNT analysis are not increasing enormously, showing that the constraining power is preserved across all bands. The constraints on $S_{8}$ remain also consistent across all $k$ bins (not shown). On the other hand, the best fit posterior of $A_{\mathrm{TA}}$ strongly depends on the $k$ bin. There is an increasing systematic shifts from the fiducial value of the TATT model ($A_{\mathrm{TA}}=1.32$) for the low $k$ bin, towards that of the NLA model fit to the same simulated galaxy sample ($A_{\mathrm{TA}}=1.71$). This tension in $A_{\mathrm{TA}}$ between the $(0.033,0.33)$ and $(0.33,3.3)$ bands is much larger than $5\sigma$. This result demonstrates a clear scale-dependent inconsistency: the IA model used in the sampling pipeline fails to capture the scale dependence intentionally introduced in the fiducial – exactly the discrepancy this test is designed to expose. This behaviour is not shared by the noBNT analysis.
Figure 9: The effective tidal alignment amplitude $A_{\mathrm{TA,eff}}(k,z)$ (Eq. II.2) constraints of all three physical scale slices derived from the contour of the Figure 8, for the constraints derived from BNT-transformed data vector at five different redshifts: z=0.0, 0.5, 1.0, 1.5, 2.0z=0.0,\;0.5,\;1.0,\;1.5,\;2.0. The errorbar on $k$-axis indicates the width of the $k$-band used in that data point, and the shaded region indicates the absolute maximum of $k$ we can explore due to our maximum $\ell_{\mathrm{max}}=5000$ choice.

### v2

@@ Abstract
[…]
@@ I Introduction
Weak gravitational lensing is fundamentally a two-dimensional probe: for a given source redshift, it measures anisotropies in the mass distribution projected along the line of sight from the sources to the observer (for a review on WL, see [4]). This makes it challenging to identify biases in the nonlinear three-dimensional modeling of the matter power spectrum, because only two-dimensional projections are observed. Consequently, the usual approach is to apply scale cuts to the WL data vector, at the cost of significant information loss. Recent stage III surveys have shown quantitatively how scale cuts limit the precision of cosmological inference [5, 6, 7].
However, [8] showed that a certain linear combination of weak-lensing observables can recover some of the three-dimensional information. This is possible because, for a lens at a given redshift, the redshift scaling of the WL effect is purely geometric and independent of the matter distribution—a property that underlies nulling techniques [9, 10]. The specific linear combination that achieves this, developed in [8], is called the Bernardeau–Nishimichi–Taruya (BNT) transform. In our previous work [11] (Paper I), we adapted and tested this transform as a targeted solution for mitigating nonlinear systematics.
We showed that a BNT-transformed data vector, combined with a well-defined angular scale filtering ($\ell$-cut in harmonic space), can effectively operate as a $k$-cut, enabling more robust handling of uncertainties in the nonlinear modeling of the three-dimensional matter power spectrum. However, in Paper I, we did not account for intrinsic alignment (IA), a known contaminant of the WL signal that can mimic lensing by correlating galaxy shapes over large distances (for reviews on IA, see [12, 13]). In this paper, we revisit a number of results obtained in paper I and reevaluate the performance of the BNT transform when IA is taken into account.
@@ II.1 Cosmic Shear
We consider a cosmological survey with measured photometric redshifts for galaxies, enabling tomographic binning with source-galaxy distributions $n_{i}(\chi)$ for the $i$th bin, where $\chi$ is the radial comoving distance. The two-dimensional cosmic-shear power spectrum $C^{i,j}(\ell)$ between tomographic bins $(i,j)$ is given by the line-of-sight projection of the non-linear three-dimensional matter power spectrum $P_{\rm nl}(k;z)$, as follows [14]:
@@
$W_{\gamma}^{i}(\chi)=\dfrac{3\Omega_{m}H_{0}^{2}}{2c^{2}}\int\text{d}\chi^{\prime}\dfrac{n_{i}(\chi^{\prime})}{a(\chi)}\dfrac{(\chi^{\prime}-\chi)\chi}{\chi^{\prime}},$ (2)
where $a(\chi)$ is the scale factor, $H_{0}$ is the Hubble constant today and $\Omega_{m}$ the total matter density today. We have assumed a flat Universe with zero spatial curvature, $K=0$. For non-flat geometry, the radial comoving distance $\chi$ has to be replaced by the comoving angular diameter distance $f_{K}(\chi)$ in Eq. 2:
@@ II.2 Intrinsic Alignment
In addition to the coherent distortions induced by gravitational lensing described by Eq. 1, intrinsic alignments (IA) generate additional angular correlations among galaxy shapes—both within a given tomographic bin, through local tidal interactions, and across widely separated bins, via long-range correlations. The relevant IA contributions are the intrinsic-intrinsic ”II”, which corresponds to the correlated galaxy shapes caused by their physical proximity, and the gravitational-intrinsic ”GI”, which corresponds to the correlation between the lensed galaxy shape caused by a large-scale structure and the intrinsic galaxy shape located in that structure. We also define the ”GG” term, which corresponds to the lensing signal (gravitational-gravitational) without IA contribution. The resulting observed correlation between tomographic bins $i$ and $j$ can be written as:
$C^{i,j}\equiv C^{i,j}_{\rm GG}+C^{i,j}_{\rm GI}+C^{i,j}_{\rm IG}+C^{i,j}_{\rm II},$ (3)
where the angular power spectra can be expressed as line-of-sight integrals of the three-dimensional power spectra $P_{\mathrm{GG}}$, $P_{\mathrm{GI}}$ and $P_{\mathrm{II}}$:
$C^{i,j}_{\rm GG}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{\gamma}(\chi)W^{j}_{\gamma}(\chi)P_{\mathrm{GG}}(k;z)$ (4)
$C^{i,j}_{\rm GI}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{\gamma}(\chi)W^{j}_{N}(\chi)P_{\mathrm{GI}}(k;z)$ (5)
$C^{i,j}_{\rm IG}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{N}(\chi)W^{j}_{\gamma}(\chi)P_{\mathrm{GI}}(k;z)$ (6)
$C^{i,j}_{\rm II}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{N}(\chi)W^{j}_{N}(\chi)P_{\mathrm{II}}(k;z)$ (7)
where $W^{i}_{N}(\chi)$ is the number density projection kernel:
$W^{i}_{N}(\chi)=b^{i}\frac{H(\chi)}{c}\,n_{i}(z(\chi)),$ (8)
with $b^{i}$ being the linear galaxy bias of ith tomographic bin and $H(\chi)$ is the Hubble constant at $\chi$. In this paper, in order to keep the focus on the intrinsic alignment, we set $b^{i}=1$.
These intrinsic correlations can mimic or bias the lensing signal, making IA a significant astrophysical systematic in weak-lensing analyses [15]. The GI, IG and II terms explicitly depend on the IA model.
@@
where $\rho_{\mathrm{crit}}$ is the critical density of the universe, $D(z)$ is the structure growth function, $z_{0}$ is a pivot redshift fixed to $z_{0}=0.62$ [5], and $\tilde{C}_{1}=5\times 10^{-14}\,h^{-2}\mathrm{M}_{\odot}^{-1}\mathrm{Mpc}^{3}$ is a normalisation constant.
@@
where $\langle a(z)\rangle_{i}$ is the $n(z)$-weighted average scale factor in tomographic bin $i$, and $a_{\mathrm{piv}}=0.769$ is the pivotal scale factor. Note that, unlike the $\eta_{\mathrm{TA}}$ parameterisation of NLA model (Eq.10), this NLA-z model can change sign as a function of redshift.
@@
$\gamma_{ij}^{\mathrm{IA}}=C_{\mathrm{1}}s_{ij}+C_{1\delta}(\delta s_{ij})+C_{2}\left[\sum_{k=0}^{2}s_{ik}s_{kj}-\frac{1}{3}\delta_{ij}s^{2}\right]+$... (12)
@@
where $A_{\mathrm{TT}}$ is the Tidal Torquing scaling amplitude and $b_{\mathrm{TA}}>0$ is the density weighting bias. To avoid introducing too many equations, please refer to the Section III. of [19] for the perturbation theory-based prescription to compute $P_{\mathrm{GI}}(k)$ and $P_{\mathrm{II}}(k)$ under the TATT model. It is worth mentioning that, for the TATT model, Eq. 9 cannot be applied. This is a consequence of Eq. 12 which implies that the redshift and $k$ dependence can no longer be factorized. For an effective description of the intrinsic alignment amplitude, compared to the standard NLA model (Eq. 10), we also define an effective tidal alignment amplitude $A_{\mathrm{TA,eff}}$:
$A_{\mathrm{TA,eff}}(k,z)\equiv\dfrac{D(z)}{\tilde{C}_{1}\rho_{\mathrm{crit}}\Omega_{m}}\sqrt{\dfrac{P_{\mathrm{II}}(k,z)}{P_{\mathrm{GG}}(k,z)}}$ (15)
The effective amplitude $A_{\mathrm{TA,eff}}$ can be calculated for any IA model used in this work (Eq.10 for NLA, Eq.11 for NLAz and Eq.14 for TATT), and will be used in Section IV to compare the model between each other.
@@ II.3 BNT Transform
It is worth mentioning that, for the NLA model, combining Eq. 9 and Eqs. 4-7 leads to a compact form for the BNT data vector $\hat{C}^{a,b}(\ell)$:
@@
In the rest of this work, we will use an Euclid-like mock with ten redshift bins weak lensing tomography [21]. In (mock-) observational practices, galaxies are binned by their photometric redshift estimates. Because photometric redshifts have finite uncertainty, the true redshift distribution $n(z)$ of each bin is broadened and overlaps with its neighboring bins, even if the original photometric binning is non-overlapping. We can compute the kernels directly from the redshift distributions, and Figure 2 displays the three possible kernel combinations as functions of $\chi$, alongside the effective kernel $\chi^{-2}\hat{W}^{2}_{\mathrm{eff}}(\chi)$. The BNT intrinsic-alignment kernel is less localized in redshift than the BNT lensing kernel, yet it remains more localized than the original noBNT lensing kernel.
@@ III.1 Scale Cuts
where $k_{\mathrm{min}}=0.0025$ Mpc-1 and $k_{\mathrm{max}}=100$ Mpc-1 is the entire computational domain in $k$-space. The $\hat{\mathfrak{R}}$ ratios represent the fractional difference between the angular power spectrum with $k$-cut (numerator) and no $k$-cut (denominator) for a given $\ell$. It literally captures the fractional power surviving the $k$-cut in $\ell$ space. As noted in Paper I, the BNT transform does not need to be computed with the true cosmology. Actually, searching for the true cosmology for which nulling would be exact can itself be exploited ([22, 10]).
Following Paper I, we also define a fractional difference threshold $\mathcal{T}_{\mathrm{FD}}$ to characterise how a $\ell$-cut is defined based on the tolerance region with $\hat{\mathfrak{R}}=1\pm\mathcal{T}_{\mathrm{FD}}$. The parameter $\mathcal{T}_{\mathrm{FD}}$ is set by the user, based on the bias level of the angular spectrum that can be tolerated (this value is the same for all elements of the data vector). This is in general determined by the accuracy required for a given cosmological parameter (e.g. $S_{8}$) to distinguish different cosmological scenario or independent measurements (e.g. Planck versus lensing measurements of $S_{8}$). In order to include the impact of both $k_{\mathrm{cuts}}$ and $\mathcal{T}_{\mathrm{FD}}$, we use the same continuous Boolean weight function $\mathcal{V}(k_{\mathrm{cuts}},\mathcal{T}_{\mathrm{FD}})$ as the paper I:
@@ IV.1 BNT and Intrinsic Alignment
We first examine the joint constraints on $(S_{8},\Omega_{m})$ in a setting where the fiducial nonlinear power spectrum is biased relative to the model used for sampling. As in Paper I, we compare constraints from the BNT and noBNT approaches for various choices of $k_{\rm cut}$, using AxionHMcode with non-zero intrinsic alignments for the fiducial data vector, while the sampling model uses HMcode. In Figure 4, we show the cosmological constraints for different physical scale cuts $k_{\rm cut}$, sampled over a TATT model with $b_{\mathrm{TA}}=0.0$, compared to the fiducial TATT model with $b_{\mathrm{TA}}=0.21$. Notably, this mismatch in $b_{\mathrm{TA}}$ does not introduce any significant additional bias in the cosmological constraints. The bias in $S_{8}$ and $\Omega_{m}$ appears when small-scale information is included, as shown by the shift of contours in Figure 4, which is primarily due to the impact of ultralight axions. However, as we progressively restrict the analysis to larger scales, the $S_{8}$ and $\Omega_{m}$ constraints converge toward the “true background cosmology” indicated by the cross in the contour plot. At $k_{\mathrm{cut}}=0.1\;\mathrm{Mpc}^{-1}$, the recovered cosmology aligns well with the fiducial. These results closely match the result in Paper I, where intrinsic alignment was not included, meaning that the biases seen in Figure 4 are not caused by the biased IA model, but effectively by the incorrect non-linear model of the mass power spectrum. We conclude that the performance of BNT is not degraded by the presence of intrinsic alignment. By contrast, the noBNT constraints lose most of their constraining power in this regime, a result also similar to Paper I. These results suggest that, even with good but imperfect knowledge of the IA model, the BNT-transform-based $\ell$-cut method remains effective for delivering unbiased cosmological constraints.
@@ IV.2 Biases induced by Intrinsic Alignment Models
We first investigate the impact of an incorrect IA model in the redshift space by using the NLA-z model from Table 2 for fiducial and NLA model for the sampling. The NLA-z model introduces a different redshift dependence than the standard NLA model, which illustrates the impact of choosing an incorrect IA model. The data vector is split in three tomographic redshift groups: $(i,j)$ or $(a,b)$ for noBNT or BNT, belonging to either $\{1,2,3\}$, $\{4,5,6\}$, or $\{7,8,9,10\}$. For each subset, we exclude cross-correlations between bins belonging to different groups, as explained in Section III.2. Figure 6 shows the resulting constraints on $A_{\mathrm{TA}}$ and $\eta_{\mathrm{TA}}$ derived from the posterior sampling. In both the noBNT and BNT analyses, the resulting contours display substantial inconsistencies across different redshift-bin groupings. Constraints from the first three tomographic bins deviate from those obtained using the full redshift range by more than $5\sigma$ in $\eta_{\mathrm{TA}}$ ($7.01\sigma$ for BNT and $7.03\sigma$ for noBNT using the mahalanobis distance [33] between the posterior means on the multi-dimensional parameter space). Likewise, the constraints from the last four tomographic bins differ from those of the first three by over $10\sigma$ (10.85 for BNT and 12.19 for noBNT), while still showing a tension exceeding $5\sigma$ (7.16 for BNT and 8.36 for noBNT) relative to the unsliced case. These findings demonstrate that strong internal inconsistencies can arise when the IA model is incorrect. Redshift grouping therefore provides a useful diagnostic for testing the validity of the assumed IA model.
@@
Figure 7: The effective tidal alignment amplitude $A_{\mathrm{TA,eff}}(z)$ (Eq. 15) constraints of all three redshift slices derived from the contour of the Figure 6, for both BNT and noBNT cases. The z-values and their errorbars of the data points are derived from the weighted mean and weighted standard deviation, when all galaxies within the $z$-slice (like $\mathbb{Z}_{1,2,3}$) are categorized to be within one single redshift distribution. For the illustration purpose only, the z-value of the noBNT and BNT data points are each shifted by $\pm 0.01$ on the plot, with their $A_{\mathrm{TA,eff}}(z)$ kept to its original constraint value.
@@
Based on the BNT and noBNT constraints shown in Figure 6, we fit an effective tidal alignment amplitude, $A_{\mathrm{TA,eff}}(z)$ (using the definition from Eq.15) for each redshift grouping. Figure 7 presents the resulting $A_{\mathrm{TA,eff}}(z)$ compared to the input NLA-z Fiducial model. The centre values and bin width along the $z$-axis are obtained from the $n(z)$-weighted mean and standard deviation, by combining the redshift slices $\mathbb{Z}_{1,2,3}$, $\mathbb{Z}_{4,5,6}$, and $\mathbb{Z}_{7,8,9,10}$ into three effective single redshift distributions. This demonstrates that, for an intrinsic alignment model with explicit redshift dependence, such an approach can serve as a viable method to constrain the detailed redshift evolution of the IA model if the sampling model cannot recover it. For instance, $A_{\mathrm{IA,eff}}(z)$ (Eq. 15) cannot cross zero with our sampling IA model (Eq. 10), but it can do so with the fiducial IA model (Eq. 11), therefore creating a discrepancy.
@@
Similarly, we investigate the impact of $k$-band cuts when the fiducial and sampling IA models differ. In this test, we adopt the $k$-dependent nonlinear TATT model, which introduces additional clustering power at large $k$-scales compared to the conventional NLA model. The fiducial data vector is generated using the TATT model with parameters fitted from the Samuroff et al. [25] study, while sampling is performed with the standard NLA model. Figure 8 shows the constraints on ($\eta_{\rm TA},A_{\rm TA}$) for three $k$-bands: $k_{\mathrm{cuts}}=(0.033,0.33)$, $(0.1,1.0)$, and $(0.33,3.3)\;\mathrm{Mpc}^{-1}$ . Notably, the noBNT analysis shows a substantial loss of constraining power, particularly for the intrinsic-alignment parameter $A_{\mathrm{TA}}$. In contrast, the contours of the BNT analysis are not increasing enormously, showing that the constraining power is preserved across all bands. The constraints on $S_{8}$ remain also consistent across all $k$ bins (not shown). On the other hand, the best fit posterior of $A_{\mathrm{TA}}$ strongly depends on the $k$ bin. There is an increasing systematic shifts from the fiducial value of the TATT model ($A_{\mathrm{TA}}=1.32$) for the low $k$ bin, towards that of the NLA model fit to the same simulated galaxy sample ($A_{\mathrm{TA}}=1.71$). This tension in $A_{\mathrm{TA}}$ between the $(0.033,0.33)$ and $(0.33,3.3)$ bands is $5.13\sigma$. This result demonstrates a clear scale-dependent inconsistency: the IA model used in the sampling pipeline fails to capture the scale dependence intentionally introduced in the fiducial – exactly the discrepancy this test is designed to expose. This behaviour is not shared by the noBNT analysis.
Figure 9: The effective tidal alignment amplitude $A_{\mathrm{TA,eff}}(k,z)$ (Eq. 15) constraints of all three physical scale slices derived from the contour of the Figure 8, for the constraints derived from BNT-transformed data vector at five different redshifts: z=0.0, 0.5, 1.0, 1.5, 2.0z=0.0,\;0.5,\;1.0,\;1.5,\;2.0. The errorbar on $k$-axis indicates the width of the $k$-band used in that data point, and the shaded region indicates the absolute maximum of $k$ we can explore due to our maximum $\ell_{\mathrm{max}}=5000$ choice.